OpenStudy (anonymous):

write the equation in vertex form y=x^2+14x+50

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

u acnd to it by completing the squares

OpenStudy (anonymous):

y=x^2+14x+50 y=x^2+2(7)x +49 -49 +50 y= (x+7)^2 +1 or (y-1)=(x+7)^2 vertex being -7,1